OpenStudy (anonymous):

The half life period of decomposition of a compound is 50 minutes. If the initial concentration is halved, the half life period is reduced to 25 minutes. What is the order of reaction?

OpenStudy (anonymous):

i think it is 1st

OpenStudy (aaronq):

It's zeroth order. Look at the equation for a zeroth order half life: \(t_{1/2}=\dfrac{[A_0]}{2k}\) assume \([A_0]\)=1, and k=1, (I'm keeping values simple to make calculations easy) then, \(t_{1/2}=\dfrac{[1]}{2(1)}=\dfrac{1}{2}\) After the first half life the initial concentration decreased by half, \([A_0]=\dfrac{1}{2}\) plugging it into the half life formula: \(t_{1/2}=\dfrac{[\dfrac{1}{2}]}{2(1)}=\dfrac{1}{4}\) The half life (\(t_{1/2}\)) was reduced by half after the first half-life period.